OpenStudy (anonymous):

Can you evaluate on the TI-84 Plus Radical Notation?

OpenStudy (eyust707):

of course

OpenStudy (eyust707):

if you ever get stuck and cannot find the function in the catalog just use a fractional exponent \[\sqrt{x} = x ^{1 \over 2} \] \[\sqrt[3]{x} = x ^ {1 \over 3} \] \[\sqrt[4]{x} = x ^ {1 \over 4} \]